### v1

# Pontus-Mpemba effects

Abstract
Mpemba effects occur after a sudden quench of control parameters if for “far” (or “hot”) initial states with respect to a final target state, the relaxation time toward the target state is shorter than for “close” (or “cold”) initial states. Following a strategy of fishermen in Pontus described by Aristotle, we introduce the Pontus-Mpemba effect as a two-step protocol which includes the time needed for preparing the system in the “far” initial state. For Markovian (classical or quantum) systems coupled to multiple reservoirs, all possible Pontus-Mpemba effects fall into three classes. The theory is applied to open two-state systems where all three classes can be realized.
Introduction.—Classical [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and quantum [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] versions of the celebrated Mpemba effect (ME) are presently attracting a lot of attention. This interest is mainly driven by the intriguing physics behind the ME, by recent trapped ion experiments [34, 35, 36], and by the promise of speeding up state preparation and manipulation protocols. For recent reviews, see Refs. [37, 38]. According to the standard definition of the ME employed in the works cited above, one considers a protocol where a set of control parameters is suddenly quenched from an initial value (defining the initial system state) to a final value (defining the target state). Comparing two different initial control parameter sets, the ME occurs if the relaxation times toward the target state are ordered in a “counterintuitive” way. For example, if one considers the relaxation times $\tau_{c,h}$ for systems starting at the respective temperatures $T_{c}<T_{h}$ and relaxing toward a target state with temperature $T_{f}<T_{c}$, the ME happens for $\tau_{h}<\tau_{c}$.
However, a different and arguably more stringent definition requires that starting from a common initial configuration, the entire two-step warming-and-freezing process (say, $T_{c}\to T_{h}\to T_{f}$) must be faster than the single-step process of freezing ($T_{c}\to T_{f}$). We call the ME in such a two-step protocol “Pontus-Mpemba effect” (PME) since this principle was already noted by Aristotle in his description of fishing strategies in the ancient town of Pontus: “Hence many people, when they want to cool hot water quickly, begin by putting it in the sun. So the inhabitants of Pontus when they encamp on the ice to fish (they cut a hole in the ice and then fish), pour warm water round their reeds that it may freeze the quicker, for they use the ice like lead to fix the reeds [39].” Indeed, if the warm-up step $T_{c}\to T_{h}$ requires more time than is saved during the subsequent freezing process $T_{h}\to T_{f}$, the Pontus fishermen could take no advantage by doing so as compared to the direct freezing route $T_{c}\to T_{f}$.
Formally, we define the PME as follows. We consider two system copies prepared in the same initial state denoted by ${\bf S}$. At time $t=0$, the first copy is subjected to a control parameter quench driving the system toward the target state ${\bf F}$ in a time span $t_{\rm{SF}}$. The second copy is first subjected to a different parameter quench, corresponding to an auxiliary environment driving the system toward the stationary state ${\bf A}$ within a time $t_{\rm{SA}}$. During the evolution toward ${\bf A}$, at a time $t_{\rm{SI}}<t_{\rm SA}$ where the intermediate state ${\bf I}$ has been reached, the system is decoupled from the auxiliary environment. By a second parameter quench, the system is then connected to the same environment used for the first system copy. The system will then reach the target state ${\bf F}$ in a time $t_{\rm{IF}}$. The PME occurs if the condition $t_{\rm{SI}}+t_{\rm{IF}}<t_{\rm{SF}}$ is satisfied. We illustrate this two-step PME protocol in Fig. 1. In the above example by Aristotle, ${\bf S}$ corresponds to the cold water, ${\bf F}$ to ice, ${\bf A}$ is induced by the sun (or by fire) and would eventually lead to evaporated water, and ${\bf I}$ is the warm water. While for the standard ME, ${\bf I}$ is a stationary (and usually thermal) state, for the PME, there is no such constraint. Indeed, ${\bf I}$ is now just a generic state along the time evolution from ${\bf S}$ to ${\bf A}$.
In order to quantify the PME, we need a proper monitoring function to extract relaxation times and distinguish different classes of PMEs, similar to what is done in the standard ME [2, 20]. The monitoring function measures the distance of the time-dependent state from the target state and has to be a monotonically non-increasing, continuous, and convex function of time [2]. While our definition of the PME and many of our arguments below are generally valid, we here focus on Markovian quantum systems [40], using the trace distance [41] as monitoring function which satisfies the above criteria. However, other choices for the monitoring function are possible [20] and lead to the same PME classification. We note that since we allow for multiple reservoirs coupled to the system, the stationary states can be current-carrying nonequilibrium steady states (NESSs). Moreover, our approach also applies to the classical limit. We introduce the concept of a “velocity field” to visualize how and when PMEs occur. Our analysis shows that all possible PMEs in Markovian systems fall into three separate classes. The theory is illustrated for two-state quantum systems, where we demonstrate the emergence of all three PME classes.
Figure 1: Velocity field $\dot{\rho}$ for the Markovian dynamics of a two-level system constrained to the $r_{1}$-$r_{2}$ plane ($r_{3}=0$) of the Bloch vector ${\bf r}(t)=(r_{1},r_{2},r_{3})^{T}$ with $|{\bf r}|\leq 1$. Here $\dot{\rho}$ can be represented by the vector field ${\bf v}({\bf r})=\dot{\bf r}$, where $\dot{\bf r}(t)$ depends only on ${\bf r}(t)$, see Eq. (1). Arrows and colors represent the direction and amplitude $(v=|{\bf v}|)$ of ${\bf v}({\bf r})$, respectively. In panel (a) [panel (b)], the steady state reached at long times is F [A], with the corresponding parameters in Eq. (1) specified in the End Matter section. The PME protocol compares the direct process ${\bf S}\to{\bf F}$ along trajectory $\Gamma_{\rm SF}$ [blue curve in panel (a)] to the two-step process composed of (i) $\bf{S}\to{\bf I}$ along trajectory $\Gamma_{\rm SA}$ [green curve in panel (b)] and (ii) ${\bf I}\to{\bf F}$ along $\Gamma_{\rm IF}$ [yellow curve in panel (a)]. If the two-step process is faster, the PME occurs.
Markovian dynamics.—The first standard form of the Lindblad equation, which is the most general generator of Markovian dynamics, is given by ($\hbar=1$)
$\frac{d\rho}{dt}=-i[H,\rho]+\sum_{m,n=1}^{N^{2}-1}C_{mn}\left(F_{m}\rho F_{n}^{\dagger}-\frac{1}{2}\left\{F_{n}^{\dagger}F_{m},\rho\right\}\right),$ (1)
[…]
$\frac{d\rho}{dt}=-i[H,\rho]+\sum_{j=1}^{N^{2}-1}\gamma_{j}\left(L_{j}\rho L_{j}^{\dagger}-\frac{1}{2}\left\{L_{j}^{\dagger}L_{j},\rho\right\}\right),$ (2)
where the jump operators $L_{j}$ are appropriate linear combinations of the operators $F_{m}$ and the rates $\gamma_{j}\geq 0$ are eigenvalues of $C$. The diagonal form (2) uses the minimal number of jump operators necessary to describe the dissipative dynamics. (Eq. (2) with a sum over more than $N^{2}-1$ terms is not in diagonal form but in first standard form under disguise.) Since $C$ in Eq. (1) has dimension $(N^{2}-1)\times(N^{2}-1)$, there are up to $(N^{2}-1)^{2}$ independent dissipator parameters. Indeed, when resorting to the diagonal form (2), the freedom lies not just in the eigenvalues but also in the basis diagonalizing $C$. We note that the classical limit follows for $H\to 0$. Similarly, the system state can be expressed by $N^{2}-1$ parameters constrained by the positivity of $\rho$. We refer to the convex space spanned by all possible $\rho$ as the $\rho$-hyperplane. For a two-level system ($N=2$), it corresponds to the Bloch unit ball. We employ the trace distance [41],
$\mathcal{D}_{T}(\rho_{1},\rho_{2})=\frac{1}{2}\mathrm{Tr}\left|\rho_{1}-\rho_{2}\right|,$ (3)
to measure the distance between two density matrices $\rho_{1}$ and $\rho_{2}$ in the $\rho$-hyperplane.
Velocity field.—The classical thermal ME can be linked to geometric properties of a free energy landscape [2]. However, this connection is not useful here since for a solution of Eqs. (1) or (2), no energy landscape needs to be computed and NESS solutions generally do not realize minima of a free energy. We instead introduce a velocity field landscape, which allows for an intuitive interpretation of the PME and which directly depends on the jump operators and the corresponding rates in Eq. (2). For Markovian systems, the relaxation dynamics in the $\rho$-hyperplane is represented in terms of a velocity field by attaching to each $\rho(t)$ the corresponding vector $\dot{\rho}(t)$. Note that one may vectorize the density matrix using the Choi isomorphism [44]. For given $H$ and $\{L_{j}\}$ in Eq. (2), we then observe that $\dot{\rho}(t)$ is a function of $\rho(t)$ only. In Fig. 1, we show the velocity field profile in the $\rho$-hyperplane for a two-state system, where we choose parameters such that all Bloch vectors ${\bf r}(t)$ have $r_{3}(t)=0$ (see End Matter). For more complex systems, in particular for $N>2$, one can visualize the velocity field in a hyperplane containing the points ${\bf S},{\bf F},$ and ${\bf A}$ (or I).
For the two panels in Fig. 1, we choose the jump operators and rates (or equivalently, the Kossakowski matrices) such that the corresponding steady states ${\bf F}$ and ${\bf A}$ with $\dot{\rho}=0$ are realized [45, 46]. Both cases correspond to different baths coupled to the system, and thus to different velocity field profiles. In Fig. 1(a), starting from the initial state S, the system evolves along a trajectory $\Gamma_{\rm SF}$ to the fixed point F, representing the first system copy in our PME definition. Note that at each point, the velocity field is tangential to the trajectory. Comparing the trajectories with starting points S and I in Fig. 1(a), we encounter a slower time evolution along $\Gamma_{\rm SF}$ as compared to $\Gamma_{\rm IF}$ since a low-velocity region is traversed. This observation implies the standard ME. Both time evolutions can be monitored by $\mathcal{D}_{T}(\rho(t),\rho_{\rm F})$, see Eq. (3), with $\rho(t)$ obtained by solving the Lindblad equation with $\rho(0)=\rho_{\rm S}$ and $\rho(0)=\rho_{\rm I}$, respectively. The trace distance will then smoothly decay to zero. Since the state F is reached only for $t\to\infty$, we impose a finite but small dimensionless cutoff $\epsilon>0$ such that for $\mathcal{D}_{T}(\rho(t_{c}),\rho_{\rm F})=\epsilon$, convergence is declared for $t>t_{c}$. Importantly, for sufficiently small $\epsilon$, the PME classification below is independent of the precise value of $\epsilon.$
The velocity field $\dot{\rho}$ offers an intuitive visualization and understanding of the PME. Following the above definition of the PME, we compare the velocity fields for two different Lindblad baths, namely the one corresponding to the target state F, see Fig. 1(a), and the one for the auxiliary state A, see Fig. 1(b). For both system copies, we start from the initial state S. The first copy evolves in the velocity field of the target state only, see the blue curve in Fig. 1(a), and thus has to propagate through a low-velocity (“slow”) region (corresponding to red color). The second copy instead first evolves in the velocity field of the auxiliary bath toward A, see the green curve in Fig. 1(b), where the propagation speed is much faster and the slow region in Fig. 1(a) is circumvented. Since one wants to approach the target state F, however, at some intermediate state I along the trajectory $\Gamma_{\rm SA}$, which is reached at time $t_{\rm SI}<t_{\rm SA}$, the system is decoupled from the auxiliary bath and connected to the target bath by performing another parameter quench. The system then evolves along the trajectory $\Gamma_{\rm IF}$, see the yellow curve in Fig. 1(a). In fact, by avoiding the slow region in Fig. 1(a), it reaches F before the first copy and the PME occurs.
Figure 2: Trace distance ${\cal D}_{T}(\rho^{({\rm F}\,({\rm A}))}(t),\rho_{\rm F})$ vs time $t$ using different target (F) and auxiliary (A) states as environment for Markovian two-state systems. As in Fig. 1, blue (green) curves show the time evolution of ${\cal D}_{T}(\rho^{({\rm F}\,({\rm A}))}(t),\rho_{\rm F})$ along ${\bf S}\to{\bf F}({\bf A})$ under the influence of the respective environment, and yellow curves show ${\cal D}_{T}(\rho^{({\rm F})}(t),\rho_{\rm F})$ along ${\bf I}\to{\bf F}$ for different intermediate states I. Dotted vertical and horizontal lines are guides to the eyes only. The parameter values taken in different panels are specified in the End Matter section. (a): Example where ${\cal D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ has a minimum and a crossing point with ${\cal D}_{T}(\rho^{({\rm F})}(t),\rho_{\rm F})$. For three states I, yellow curves show the trace distance along $\Gamma_{\rm IF}$. These three cases realize all different PME types discussed in the main text. (b): Example where ${\cal D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ has a minimum but no crossing point with ${\cal D}_{T}(\rho^{({\rm F})}(t),\rho_{\rm F})$. (c): Example for the case in Eq. (6), where no minimum in ${\cal D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ exists.
PME classification.—We turn to a classification of possible PME types in Markovian systems. To begin, let us analyze the relative position between the trajectories $\Gamma_{\rm SF}$ and $\Gamma_{\rm SA}$. A first case arises if, starting from ${\bf S}$, the system dynamics under the auxiliary bath (indicated by the superscript (A) below) initially reduces the trace distance to both states A and F, while at later times the trajectory moves away from F. This first case is therefore characterized by the two conditions
$\partial_{t}\mathcal{D}_{T}(\rho^{({\rm A})}(t=0),\rho_{\rm F})<0,$
$\exists\>t_{M}>0\>\mid\>\partial_{t}\mathcal{D}_{T}(\rho^{({\rm A})}(t=t_{M}+0^{+}),\rho_{\rm F})>0.$ (4)
The second condition means that the trace distance $\mathcal{D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ has a minimum at $t=t_{M}$ corresponding to the state along $\Gamma_{\rm SA}$ closest to F. If there are several minima, $t_{M}$ refers to the first minimum along $\Gamma_{\rm SA}$. We further distinguish two sub-cases according to the conditions
$\partial_{t}\mathcal{D}_{T}(\rho^{({\rm A})}(t=0),\rho_{\rm F})\lessgtr\partial_{t}\mathcal{D}_{T}(\rho^{({\rm F})}(t=0),\rho_{\rm F}),$ (5)
which are illustrated in Figs. 2(a) and 2(b), respectively. These two sub-cases are characterized by the existence, or not, of a finite-time crossing point between the trace distance curves $\mathcal{D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ and $\mathcal{D}_{T}(\rho^{({\rm F})}(t),\rho_{\rm F})$, where the superscript (F) means that the dynamics is driven by the target environment.
A second case arises if $\mathcal{D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ monotonically decreases in time. It is similar to the first case but the minimum along $\Gamma_{\rm SA}$ in the trace distance to F is now washed out, i.e., $t_{M}\to\infty$. We do not discuss this case further since the resulting PME behavior is already covered by the first case. Finally, a third case exists if, starting from S, the system moves away from the target state F, thus increasing the trace distance from it, while being attracted to A. One then arrives at the condition
$\partial_{t}\mathcal{D}_{T}(\rho^{({\rm A})}(t=0),\rho_{\rm F})>0.$ (6)
An example for this case is illustrated in Fig. 2(c). Finally, if quantum effects due to $H$ dominate over the dissipative dynamics in Eq. (1), spiral trajectories may emerge in the $\rho$-hyperplane, implying damped oscillations of $\mathcal{D}_{T}(\rho^{({\rm A})}(t),\rho_{\rm F})$ (see End Matter).
We now provide the general PME classification. For the second system copy (subjected to the two-step protocol), the intermediate state I is reached at time $t_{\rm SI}$ along the trajectory $\Gamma_{\rm SA}$. One now performs a parameter quench to decouple the system from the auxiliary bath and instead couples it to the target bath, see Fig. 1,
$\rho(t)=\begin{cases}\rho^{({\rm A})}(t),&t<t_{\rm SI},\\ \rho^{({\rm F})}(t),&t\geq t_{\rm SI}.\end{cases}$ (7)
We can then distinguish three different PME types depending on the location of the corresponding intermediate state I. If the condition
$\mathcal{D}_{T}(\rho^{({\rm A})}(t_{\rm SI}),\rho_{\rm F})<\mathcal{D}_{T}(\rho^{({\rm F})}(t_{\rm SI}),\rho_{\rm F})$ (8)
holds, we define the weak type-I PME. An example for this case is given by the lowest-lying yellow curve in Fig. 2(a). It is characterized by the absence of a crossing point between the trace distances for the dynamics of the first and the second system copy, with the state of the second copy being closer to ${\bf F}$ at all times. On the other hand, for
$\mathcal{D}_{T}(\rho_{\rm S},\rho_{\rm F})>\mathcal{D}_{T}(\rho^{({\rm A})}(t_{\rm SI}),\rho_{\rm F})>\mathcal{D}_{T}(\rho^{({\rm F)}}(t_{\rm SI}),\rho_{\rm F}),$ (9)
we speak of a weak type-II PME. Examples are given by the middle yellow curve in Fig. 2(a) and by the lower yellow curve in Fig. 2(b), with a finite-time crossing point between the trace distances for both system copies.
Finally, the most elusive case, dubbed strong PME, may arise for the third case discussed above, see Eq. (6). It is realized if
$\mathcal{D}_{T}(\rho^{({\rm A})}(t_{\rm SI}),\rho_{\rm F})>\mathcal{D}_{T}(\rho_{\rm S},\rho_{\rm F}).$ (10)
Examples for the strong PME are given by the upper yellow curves in Figs. 2(a,b) and by the yellow curve in Fig. 2(c). This is the hardest and most counterintuitive PME to achieve, but it is closest in spirit to the Pontus example by Aristotle [39]. From an application point of view, however, we expect that the two weak PME cases will provide faster relaxation times toward F in practice.
Two-level systems.—As probably simplest application, we now turn to Markovian two-level systems. The results in Figs. 1 and 2 have been obtained for this case and show that all three PME classes can already be realized for $N=2$. For technical details and the parameter values in Figs. 1 and 2, see the End Matter section. For $N=2$, we express the traceless operators $F_{n}$ in the Lindblad equation Eq. (1) by the three Pauli matrices $\sigma_{n=1,2,3}$, and the system state is $\rho(t)=\frac{1}{2}(\mathbb{I}+\sum_{n}r_{n}(t)\sigma_{n})$ with the Bloch vector ${\bf r}(t)$. The trace distance (3) between two states $\rho_{1}$ and $\rho_{2}$ with Bloch vectors ${\bf r}_{1}$ and ${\bf r}_{2}$ is then given by ${\cal D}_{T}(\rho_{1},\rho_{2})=\frac{1}{2}|{\bf r}_{1}-{\bf r}_{2}|$. Furthermore, in Eq. (1), the system Hamiltonian is $H=\sum_{n}h_{n}\sigma_{n}$ with a real-valued vector ${\bf h}=(h_{1},h_{2},h_{3})^{T}$, while the Kossakowski matrix has the general form $C=\left(\begin{array}[]{ccc}C_{11}&C_{12}&C_{31}^{*}\\ C_{12}^{*}&C_{22}&C_{23}\\ C_{31}&C_{23}^{*}&C_{33}\end{array}\right),$ with real-valued diagonal elements and three independent complex-valued off-diagonal entries. To visualize the system dynamics within the Bloch unit ball, the velocity field $\dot{\rho}$ can be written in Bloch representation in terms of the vector field ${\bf v}({\bf r})=\dot{{\bf r}}$, where the Lindblad equation gives
$\dot{\bf r}={\bf v}({\bf r})=2{\bf\Lambda}\cdot{\bf r}+{\bf b}.$ (11)
The $3\times 3$ matrix ${\bf\Lambda}$ and the vector ${\bf b}$ are expressed in terms of $C$ and ${\bf h}$ (see End Matter). Using these equations, we obtain Figs. 1 and 2 and the corresponding PME types. We note that for $N>2$, similar representations with ${\bf v}=\dot{\bf r}$ apply by using the $(N^{2}-1)$-dimensional coherence vector ${\bf r}$ generalizing the Bloch vector [47, 48].
As demonstrated above, all three PME classes could be realized experimentally in Markovian two-state systems. There are various physical platforms to achieve this goal with reservoir engineering techniques, e.g., trapped ions [34, 35], superconducting qubits [49], or ultracold atoms [50]. Using the specific Kossakowski matrices $C$ in the End Matter section, the jump operators $L_{j}$ in Eq. (2) follow by diagonalizing $C=U^{\dagger}\Lambda U$, where the unitary matrix $U$ contains the eigenvectors and the diagonal matrix $\Lambda$ the eigenvalues, as $L_{j}=\sum_{i}U_{ji}\sigma_{i}$. The corresponding transitions rates are given by $\gamma_{j}=\Lambda_{jj}$. The time-dependent trace distance can then be measured by quantum state tomography techniques. We conclude that an experimental realization of the PME and a test of our classification into three separate PME types seems in direct reach.
Conclusions.—We have introduced the PME as a generalized protocol realizing Mpemba physics by taking into account the time needed to prepare the initially “far” state. This protocol is more demanding but also more stringent than the standard Mpemba protocol. We find three different PME classes for Markovian systems, which have been demonstrated explicitly for two-state quantum systems. We hope that our work will motivate future experimental and theoretical work investigating and possibly extending the protocol suggested here.
Acknowledgements.
We thank D. Giuliano and X. Turkeshi for discussions and acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Projektnummer 277101999 - TRR 183 (project B02) and under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769.